OpenStudy (anonymous):

use the method of substitution to solve. 4a=2-b 4a=4-b

OpenStudy (anonymous):

doesn't seem very likely that \(4a=2-b\) but at the same time \(4a=4-b\) does it?

OpenStudy (anonymous):

how can \(2-b=4-b\) ??

OpenStudy (anonymous):

that was the problem that i was given

OpenStudy (anonymous):

ok fine, you were given the problem, but that doesn't mean there is a solution

OpenStudy (anonymous):

you might write "inconsistent" or "no solution " or something similar

OpenStudy (anonymous):

that is an open so ill choose that thanks